Short Report – QA/QC Meeting Mar 2019

Meeting Date: 28 Mar 2019


LH, MG, IB, SG, PM, IF, TB, YaL, RM, PL (10)


Talking Points


  • If you are (Tech-)SRP, please attend QA/QC meetings or tell LH if you can’t, needed for planning of the meetings.
  • SRP & Tech-SRP: please prepare short statement about your site and send to to LH (max. 2 sentences)
  • The purpose of QA/QC meetings is to check on current, incoming data. SRPs choose specific issues we should look at together and discuss in the group. Fluxes are checked if the respective SRP wishes to do so.

This meeting we discussed the following topics

  • Calibration-gas issue: the CO2 calibration gas that was used to calibrate IRGAs was off by 3%. This affects fluxes from all sites since 2016. To correct for this issue, we need correction factors for all affected time periods. Level-1 fluxes 2016-2018 will then be recalculated.
  • It is important that SRPs run on-demand flux calculations on the RDS, because the automatic flux calculations shown in the QuickViews are not working all the time. RDS fluxes can be used to fill data gaps in the QVs.
  • Post-processing questions (e.g. gap-filling) better fit in separate meetings (flux_processing meetings) because they are not directly related to checks on new and incoming data, and because they are relatively time intensive to discuss.
  • For FLUXNET data sharing, we need the eastward and northward separations of the sensors in relation to the main tower, i.e. the sensor positions as meters distance. For the future, we aim to create maps of all sites that show the location of the sensors in the field. It is currently not clear which software will be used for that.


Site Reports


SRP (SG): fluxes from 2016-2019 has been gap filled and partitioned. the computer stopped working on Saturday 23rd of March 19 and Ivo is working on installing a new one. the last computer went off last year in May when we have the data gap!
T-SRP (IB): The mast is still behaving a bit difficult, it stopped again but after a hard reset on a nice warm day it was running again. Sadly only a few days later, the computer broke down on the 23.3.19. So at the moment there is no flux data until we can fix this problem.


SRP (IF): Data gaps of QCL fluxes end of January and in the mid of February. Replacing the CF card solved the problem, but may have influenced the measurement too (fluxes closer to zero after the last gap). Installation of new soil sensors on the northern parcel on March 6th.
T-SRP: –


SRP (MG, LH): EC fluxes look OK so far, no major gaps. Some defective soil sensors need replacement once site is snow free.
T-SRP: –


SRP (LH): No major gaps in eddy data, meteo data generally OK, but replicate measurement TS_GF1_0x05_2_Avg_degC_Avg gives no data.
T-SRP: –


SRP (RM): Station is down since 26th February 2019.
T-SRP: –


SRP: –
T-SRP (TB): After a troublesome start in the year we (Ivo) replaced the Irga from the eddy system and are running now finally a LI7200 again. Besides that I started to work on the dendrometer problems. A lot were destroyed due heavy damage of trees during (end of) 2018. Growing of the wild garlic started very early this year (again).


SRP (RM): Station runs although there are some incomplete data files at the beginning of the month.
T-SRP: –


Last Updated on 31 May 2019 17:50